OpenStudy (anonymous):

the ruts of the function m^6 + 3m - 7 = 0

OpenStudy (anonymous):

Delta = b^2 - 4 ( a ) ( c ) = 2^2 - 4 ( 1 ) ( 3 ) = 4 - 12 = - 8 delta is negative so the function has no roots

OpenStudy (amistre64):

yes, they did all the work for you, and they used the proper values

OpenStudy (amistre64):

its not that it doesnt have any roots, it just has no real roots

OpenStudy (amistre64):

the square root of -8 still has solutions, just complex solutions